### v4

@@ 3.1 Deepfake Generation
BlendFace shiohara2023blendface ICCV’23 ➊➌➍➏ G. Unable to handle occlusion and extreme lighting. The identity features obtained from the de-entanglement are fed to the generator as an identity loss function, which guides the generator to generate an image to fit the source image identity information.
@@
FlowFace zeng2023flowface AAAI’23 ➊➋➌➍ G.+3DMM Altered target image lighting details. It consists of face reshaping network and face exchange network, which better solves the influence of the difference between source and target face contours on the face exchange work.
@@
S2Swap liu2023high MM’23 ➊➋➌➑ G.+3D Poor ability to preserve face feature attributes. Achieving high-fidelity face swapping through semantic disentanglement and structural enhancement.
@@
StableSwap zhu2024stableswap TMM’24 ➊➌ G.+3D Unable to handle extreme skin color differences. Utilizing a multi-stage identity injection mechanism effectively combines facial features from both the source and target to produce high-fidelity face swapping.
Difussion DiffFace kim2022diffface arXiv’22 ➊➌ D. Facial lighting attributes are altered. Claims to be the first diffusion model-based face exchange framework.
@@
DiffSwap zhao2023diffswap CVPR’23 ➊➌ D. Poor ability to handle facial occlusion. Reenvisioning face swapping as conditional inpainting to harness the power of the diffusion model.
FaceX han2023generalist arXiv’23 ➊➌➏ D. Unable to handle extreme skin color differences. A novel facial all-rounder model capable of performing various facial tasks.
@@
Liu et alliu2024towards arXiv’24 ➊➋➌ D. Poor ability to preserve face feature attributes. Conditional diffusion model introduces identity and expression encoders components, achieving a balan- ce between identity replacement and attribute preservation during the generation process.
@@
Alternative Cui et alcui2023face CVPR’23 ➊➋ Other Altered target image lighting details. Introducing a multiscale transformer network focusing on high-quality semantically aware corresponden- ces between source and target faces.
@@
TransFS cao2023transfs FG’23 ➊➋➓ Other Unable to process posture differences effectively. The identity generator is designed to reconstruct high-resolution images of specific identities, and an attention mechanism is utilized to enhance the retention of identity information.
@@
Wang et alwang2024efficient TMM’24 ➊➋ Other Poor ability to handle facial occlusion. A Global Residual Attribute-Preserving Encoder (GRAPE) is proposed, and a network flow considering the facial landmarks of the target face was introduced, achieving high-quality face swapping.
3.1.1 Face Swapping
@@
In this section, we review face swapping methods from the perspective of basic architecture, which can be mainly divided into four categories and summarized in  Table 2.
@@
$\bullet$ Traditional Graphics. As representative early implementations, traditional graphics methods in the implementation path can be divided into two categories: 1) Key information matching and fusion. Methods bitouk2008face ; sunkavalli2010multi ; nirkin2018face ground in critical information matching and fusion are geared towards substituting corresponding regions by aligning key points within facial regions of interest (ROIs), such as the mouth, eyes, nose, and mouth, between the source and target images. Following this, additional procedures such as boundary blending and lighting adjustments are executed to produce the resulting image. Bitouk et albitouk2008face accomplish automated face replacement by constructing a substantial face database to locate faces with akin poses and lighting conditions for substitution. Meanwhile, Nirkin et alnirkin2018face enhance keypoint matching and segmentation accuracy by incorporating a Fully Convolutional Network (FCN) into their method. 2) The construction of a 3D prior model for facial parameterization. Methods blanz2004exchanging ; dale2011video based on constructing a 3D prior and introducing a facial parameter model often involve building a facial parameter model using 3DMM technology based on a pre-collected face database. After matching the facial information of the source image with the constructed face model, specific modifications are made to the relevant parameters of the facial parameter model to generate a completely new face. Dale et aldale2011video utilize 3DMM to track facial expressions in two videos, enabling face swapping in videos. Some methods lin2012face ; guo2019face explore scenarios involving significant pose differences between the source and target images. Lin et allin2012face construct a 3D face model from frontal faces, renderable in any pose. Guo et alguo2019face utilize plane parameterization and affine transformation to establish a one-to-one dense mapping between 2D graphics. Traditional computer graphics methods solve basic face-swapping problems, exploring full automation to enhance generalization. However, these methods are constrained by the need for similarities in pose and lighting between source and target images. They also face challenges like low image resolution, modification of target attributes, and poor performance in extreme lighting and occlusion scenarios.
$\bullet$ Generative Adversarial Network. GAN-based methods aim to obtain realistic images generated through adversarial training between the generator and the discriminator, becoming the mainstream face swapping approach. According to different improvement objectives, methods can be classified into seven categories:
@@
1) Early GAN-based methods bao2017cvae ; moniz2018unsupervised ; liu2023deepfacelab ; zhu2020deepfakes address issues related to the similarity of pose and lighting between source and target images. DepthNets moniz2018unsupervised combines GANs with 3DMM to map the source face to any target geometry, not limited to the geometric shape of the target template. This allows it to be less affected by differences in pose between the source and target faces. However, they face challenges in generalizing the trained model to unknown faces.
@@
2) Improved Generalizability. To improve the model’s generalization, many efforts sun2018hybrid ; nirkin2019fsgan ; chen2020simswap are made to explore solutions. Combining GANs with VAEs, the model bao2018towards ; natsume2018rsgan encodes and processes different facial regions separately. FSGAN nirkin2019fsgan integrates face reenactment with face swap, designing a facial blending network to mix two faces seamlessly. SimSwap chen2020simswap introduces an identity injection module to avoid integrating identity information into the decoder. However, these methods suffer from low resolution and significant attribute loss and need help to handle facial occlusions effectively.
@@
3) Resolution Upgrading. Some methods zhu2021one ; xu2022designing ; jiang2023styleipsb provide solutions to enhance the resolution of generated images. MegaFS zhu2021one introduces the first single-lens face swapping method at the million-pixel level. The face encoder no longer compresses facial information but represents it in layers, achieving more detailed preservation. StyleIPSB jiang2023styleipsb constrains semantic attribute codes within the subspace of StyleGAN, thereby fixing certain semantic information during face swapping to preserve pore-level details.
@@
4) Geometric Detail Preservation. To capture and reproduce more facial geometric details, some methods wang2021hififace ; ren2023reinforced ; zeng2023flowface ; zhang2023flowface++ introduce 3DMM into GANs, enabling the incorporation of 3D priors. HifiFace wang2021hififace introduces a novel 3D shape-aware identity extractor, replacing traditional face recognition networks to generate identity vectors that include precise shape information. FlowFace zeng2023flowface introduces a two-stage framework based on semantic guidance to achieve shape-aware face swapping. FlowFace++ zhang2023flowface++ improves upon FlowFace by utilizing a pre-trained Mask Autoencoder to convert face images into a fine-grained representation space shared between the target and source faces. It further enhances feature fusion for both source and target by introducing a cross-attention fusion module. However, most of the aforementioned methods often struggle to effectively handle occlusion issues.
@@
5) Facial Masking Artifacts. Some methods li2019faceshifter ; nirkin2019fsgan ; liu2023fine ; rosberg2023facedancer have partially alleviated the artifacts caused by facial occlusion. FSGAN nirkin2019fsgan designs a restoration network to estimate missing pixels. E4S liu2023fine redefines the face-swapping problem as a mask-exchanging problem for specific information. It utilizes a mask-guided injection module to perform face swapping in the latent space of StyleGAN. However, overall, the methods above have not thoroughly addressed the issue of artifacts in generated images under extreme occlusion conditions.
@@
6) Trade-offs between Identity-replacement and Attribute-retention. In addition to the occlusion issues that need further handling, researchers gao2021information ; shiohara2023blendface discover that the balance between identity replacement and attribute preservation in generated images seems akin to a seesaw. Many methods xu2022styleswap ; kim2022smooth ; liu2023high explore the equilibrium between identity replacement and attribute retention. InfoSwap gao2021information aims to decouple identity and attribute information in face swapping by leveraging the information bottleneck theory. It seeks controlled swapping of identity between source and target faces. StyleSwap xu2022styleswap introduces a novel swapping guidance strategy, the ID reversal, to enhance the similarity of facial identity in the output. Shiohara et alshiohara2023blendface propose BlendFace, using an identity encoder that extracts identity features from the source image and uses it as identity distance loss, guiding the generator to produce facial exchange results.
@@
7) Model Light-weighting is also an important topic with profound implications for the widespread application of models. FastSwap yoo2023fastswap achieves this by innovating a decoder block called Triple Adaptive Normalization (TAN), effectively integrating identity information from the source image and pose information from the target image. XimSwap ancilotto2023ximswap modifies the design of convolutional blocks and the identity injection mechanism, successfully deploying on STM32H743.
@@
Table 3: Overview of representative face reenactment methods. Notations: ➊ Voxceleb, ➋ Self-build, ➌ Voxceleb2, ➍ TalkingHead-1KH, ➎ CelebV-HQ, ➏ VFHQ, ➐ RaFD, ➑ VGGFace, ➒ CelebV, ➓FFHQ. Abbreviations: Expression (Exp).
@@
Face2Face thies2016face2face CVPR’16 Exp, Pose Using 3D facial reconstruction, transfer facial expressions and pose from a driving character to a target character via affine transforma -tion, then generate the final video through rendering techniques.
Kim et alkim2018deep TOG’18 Exp, Pose, Blink Using synthesized rendering images of a parameterized face model as input, creating lifelike video frames for the target actor.
Kim et alkim2019neural TOG’19 Lip,Exp, Pose Built on a recurrent generative adversarial network, it employs a hierarchical neural face renderer to synthesize realistic video frames.
Head2Head koujan2020head2head FG’21 Exp, Pose, Blink, Gaze The model consists of two stages: facial reconstruction and tracking in the first stage, followed by video rendering in the second stage.
HeadGAN doukas2021headgan ECCV’21 Exp, Pose Using 3DMM for facial modeling provides a 3D prior to the GAN, effectively guiding the generator to accurately recover pose and expression from the target frame.
Face2Faceρ yang2022face2face ECCV’22 Exp, Pose Decoupling the actor’s facial appearance and motion information with two separate encodings allows the network to learn facial app -earance and motion priors.
PECHead gao2023high CVPR’23 Lip, Exp, Pose ➌➍➎➏ A novel multi-scale feature alignment module for motion perception is proposed to minimize distortion during motion transmission.
@@
X2Face wiles2018x2face CVPR’18 Exp, Pose The training is realized in two stages: the first stage guides the generation frames towards the driving frames and the second stage accomplishes the preservation of the source identity information through the identity loss function.
Zakharov et alzakharov2019few ICCV’19 Exp, Pose ➊➌ Proposed a meta-learning framework for adversarial generative models, reducing the required training data size.
FReeNet zhang2020freenet CVPR’20 Exp, Pose moore2010multi A new triple perceptual loss is proposed to richly reproduce facial details of the face.
Zakharov et alzakharov2020fast ECCV’20 Exp, Pose Decomposed facial information into two layers for modeling: the first layer synthesizes coarse images related to the pose using a small neural network, and the second layer defines texture images unrelated to the pose, containing high-frequency details.
MarioNETte ha2020marionette AAAI’20 Exp, Pose ➊➒ Introduced image attention blocks, target feature alignment modules, and landmark transformers, enhancing the model’s performance when generalizing to unknown individual identities.
DG hsu2022dual CVPR’22 Exp, Pose ➊➌➐ A proposed dual generator model network for large pose face reproduction.
Doukas et aldoukas2023free TPAMI’23 Exp, Pose, Gaze kellnhofer2019gaze360 zhang2015appearance Eye gaze control in the generated video is implemented to further enhance visual realism.
MetaPortrait zhang2023metaportrait CVPR’23 Exp, Pose By establishing dense facial keypoint matching, accurate deformation field prediction is achieved, and the model training is expedited based on the meta-learning philosophy.
Yang et alyang2023learning AAAI’24 Exp, Pose ➊➌➍ The facial tri-plane is represented by canonical tri-plane, identity deformation, and motion components, achieving face reenactment without the need for 3D parameter model priors.
FSRTrochow2024fsrt CVPR’24 Exp, Pose The Transformer-based encoder-decoder effectively encodes attributes and improves action transmission quality.
@@
HiDe-NeRF li2023one CVPR’23 Lip, Exp, Pose ➊➌➍ High-fidelity and free-viewing talking head synthesis using deformable neural radiation fields.
HyperReenact bounareli2023hyperreenact ICCV’23 Exp, Pose ➊➌ Exploiting the effectiveness of hypernetworks in real image inversion tasks and extending them to real image manipulation.
Stylemask bounareli2023stylemask FG’23 Exp, Pose This work optimizes a Mask Network and combines it with StyleGAN2’s style potential space S in order to achieve the separation of facial pose and expression of the target image from the identity features of the source image.
Bounareli et albounareli2024one IJCV’24 Exp, Pose ➊➌ In GANs’ latent space, head pose and expression changes are decoupled, achieving near-real outputs through real image embedding.
@@
ICface tripathy2020icface WACV’20 Exp, Pose The model is decoupled and driven by interpretable control signals that can be obtained from multiple sources such as external driving videos and manual controls.
@@
Oorloff et aloorloff2023robust ICCV’23 Lip, Exp, Pose Identity and attribute decomposition are realized in StyleGAN2’s latent space, and a cyclic manifold adjustment technique enhances facial reconstruction results.
@@
Xue et al. xue2023high TOMM’23 Exp, Pose ➊➌ High-fidelity facial generation is achieved by using information-rich Projected Normalized Coordinate Code (PNCC) and eye maps, replacing sparse facial landmark representations.
$\bullet$ Diffusion-based. The latest studies kim2022diffface ; zhao2023diffswap ; han2023generalist ; liu2024towards in this area produce promising generation results. DiffSwaps zhao2023diffswap redefines the face swapping problem as a conditional inpainting task. Liu et alliu2024towards introduce a multi-modal face generation framework and achieved this by introducing components such as balanced identity and expression encoders to the conditional diffusion model, striking a balance between identity replacement and attribute preservation during the generation process. As a novel facial generalist model, FaceX han2023generalist can achieve various facial tasks, including face swapping and editing. Leveraging the pre-trained StableDiffusion blattmann2023stable has significantly improved the quality and model training speed.
$\bullet$ Alternative Techniques. Some methods stand independently from the above classifications that are discussed here collectively. Fast Face-swap korshunova2017fast views the identity swap task as a style transfer task, achieving its goals based on VGG-Net. However, this method has poor generalization. Some methods cui2023face ; cao2023transfs apply the Transformer architecture to face swapping tasks. Leveraging a facial encoder based on the Swin Transformer liu2021swin , TransFS cao2023transfs obtains rich facial features, enabling facial swapping in high-resolution images. ReliableSwap yuan2023reliableswap enhances the model’s identity preservation capabilities by constructing a reliable supervisor called the "cyclic triplet." However, it has limitations in preserving attribute information.
3.1.2 Face Reenactment

### v5

@@ 3.1. Deepfake Generation
BlendFace (shiohara2023blendface, ) ICCV’23 ➊➌➍➏ G. Unable to handle occlusion and extreme lighting. The identity features obtained from the de-entanglement are fed to the generator as an identity loss function, which guides the generator to generate an image to fit the source image identity information.
@@
FlowFace (zeng2023flowface, ) AAAI’23 ➊➋➌➍ G.+3DMM Altered target image lighting details. It consists of face reshaping network and face exchange network, which better solves the influence of the difference between source and target face contours on the face exchange work.
@@
S2Swap (liu2023high, ) MM’23 ➊➋➌➑ G.+3D Poor ability to preserve face feature attributes. Achieving high-fidelity face swapping through semantic disentanglement and structural enhancement.
@@
StableSwap (zhu2024stableswap, ) TMM’24 ➊➌ G.+3D Unable to handle extreme skin color differences. Utilizing a multi-stage identity injection mechanism effectively combines facial features from both the source and target to produce high-fidelity face swapping.
@@
Difussion DiffSwap (zhao2023diffswap, ) CVPR’23 ➊➌ D. Poor ability to handle facial occlusion. Reenvisioning face swapping as conditional inpainting to harness the power of the diffusion model.
@@
Liu et al(liu2024towards, ) CVPR’24 ➊➋➌ D. Poor ability to preserve face feature attributes. Conditional diffusion model introduces identity and expression encoders components, achieving a balan- ce between identity replacement and attribute preservation during the generation process.
DiffFace (kim2025diffface, ) PR’25 ➊➌ D. Facial lighting attributes are altered. Claims to be the first diffusion model-based face exchange framework.
Baliah et al(baliah2025realistic, ) WACV’25 ➌➐ D. Unable to handle extreme pose and expressions. Achieve improvements in identity fidelity, pose consistency, and model generalization capability.
PixSwap (kim2025pixswap, ) WACV’25 ➋➌ D. Unable to handle extreme pose and eyes control. Achieve accurately reflecting the source identity and generating more high-quality images.
@@
Alternative Cui et al(cui2023face, ) CVPR’23 ➊➋ Other Altered target image lighting details. Introducing a multiscale transformer network focusing on high-quality semantically aware corresponden- ces between source and target faces.
@@
TransFS (cao2023transfs, ) FG’23 ➊➋➓ Other Unable to process posture differences effectively. The identity generator is designed to reconstruct high-resolution images of specific identities, and an attention mechanism is utilized to enhance the retention of identity information.
@@
Wang et al(wang2024efficient, ) TMM’24 ➊➋ Other Poor ability to handle facial occlusion. A Global Residual Attribute-Preserving Encoder (GRAPE) is proposed, and a network flow considering the facial landmarks of the target face was introduced, achieving high-quality face swapping.
CanonSwap (luo2025canonswap, ) ICCV’25 ➊➎ other Poor ability to preserve face feature attributes. Disentangles motion and appearance for video face swapping, enabling precise identity transfer with pre- served temporal dynamics .
3.1.1. Face Swapping
@@
In this section, we review face swapping methods from the perspective of basic architecture, which can be mainly divided into four categories and summarized in  Tab. 2.
@@
$\bullet$ Traditional Graphics. As representative early implementations, traditional graphics methods in the implementation path can be divided into two categories: 1) Key information matching and fusion. Methods (bitouk2008face, ; sunkavalli2010multi, ; nirkin2018face, ) ground in critical information matching and fusion are geared towards substituting corresponding regions by aligning key points within facial regions of interest (ROIs), such as the mouth, eyes, nose, and mouth, between the source and target images. Following this, additional procedures such as boundary blending and lighting adjustments are executed to produce the resulting image. Bitouk et al(bitouk2008face, ) accomplish automated face replacement by constructing a substantial face database to locate faces with akin poses and lighting conditions for substitution. Meanwhile, Nirkin et al(nirkin2018face, ) enhance keypoint matching and segmentation accuracy by incorporating a Fully Convolutional Network (FCN) into their method. 2) The construction of a 3D prior model for facial parameterization. Methods (blanz2004exchanging, ; dale2011video, ) based on constructing a 3D prior and introducing a facial parameter model often involve building a facial parameter model using 3DMM technology based on a pre-collected face database. After matching the facial information of the source image with the constructed face model, specific modifications are made to the relevant parameters of the facial parameter model to generate a completely new face. Dale et al(dale2011video, ) utilize 3DMM to track facial expressions in two videos, enabling face swapping in videos. Some methods (lin2012face, ; guo2019face, ) explore scenarios involving significant pose differences between the source and target images. Lin et al(lin2012face, ) construct a 3D face model from frontal faces, renderable in any pose. Guo et al(guo2019face, ) utilize plane parameterization and affine transformation to establish a one-to-one dense mapping between 2D graphics. Traditional computer graphics methods solve basic face-swapping problems, exploring full automation to enhance generalization. However, these methods are constrained by the need for similarities in pose and lighting between source and target images. They also face challenges like low image resolution, modification of target attributes, and poor performance in extreme lighting and occlusion scenarios.
$\bullet$ Generative Adversarial Network. GAN-based methods can be classified into seven categories:
@@
1) Early GAN-based methods (bao2017cvae, ; moniz2018unsupervised, ; liu2023deepfacelab, ; zhu2020deepfakes, ) address issues related to the similarity of pose and lighting between source and target images. DepthNets (moniz2018unsupervised, ) combines GANs with 3DMM to map the source face to any target geometry, not limited to the geometric shape of the target template. This allows it to be less affected by differences in pose between the source and target faces. However, they face challenges in generalizing the trained model to unknown faces.
@@
2) Improved Generalizability. To improve the model’s generalization, many efforts (sun2018hybrid, ; nirkin2019fsgan, ; chen2020simswap, ) are made to explore solutions. Combining GANs with VAEs, the model (bao2018towards, ; natsume2018rsgan, ) encodes and processes different facial regions separately. FSGAN (nirkin2019fsgan, ) integrates face reenactment with face swap, designing a facial blending network to mix two faces seamlessly. SimSwap (chen2020simswap, ) introduces an identity injection module to avoid integrating identity information into the decoder. These methods remain limited by low resolution, attribute degradation, and poor handling of facial occlusions.
@@
3) Resolution Upgrading. Some methods (zhu2021one, ; xu2022designing, ; jiang2023styleipsb, ) aim to improve the resolution of generated faces. MegaFS (zhu2021one, ) introduces the first single-lens face swapping method at the million-pixel level. The encoder no longer compresses facial information but represents it in layers, achieving more detailed preservation. StyleIPSB (jiang2023styleipsb, ) constrains semantic attribute codes within the subspace of StyleGAN, thereby fixing semantic information during face swapping to preserve pore-level details.
@@
4) Geometric Detail Preservation. To capture and reproduce more facial geometric details, some methods (wang2021hififace, ; ren2023reinforced, ; zeng2023flowface, ; zhang2023flowface++, ) introduce 3DMM into GANs, enabling the incorporation of 3D priors. HifiFace (wang2021hififace, ) introduces a novel 3D shape-aware identity extractor, replacing traditional face recognition networks to generate identity vectors that include precise shape information. FlowFace (zeng2023flowface, ) introduces a two-stage framework based on semantic guidance to achieve shape-aware face swapping. FlowFace++ (zhang2023flowface++, ) improves upon FlowFace by utilizing a pre-trained Mask Autoencoder to convert face images into a fine-grained representation space shared between the target and source faces. It further enhances feature fusion for both source and target by introducing a cross-attention fusion module. However, most of the aforementioned methods often struggle to effectively handle occlusion issues.
@@
5) Facial Masking Artifacts. Some methods (li2019faceshifter, ; nirkin2019fsgan, ; liu2023fine, ; rosberg2023facedancer, ) have partially alleviated the artifacts caused by facial occlusion. FSGAN (nirkin2019fsgan, ) designs a restoration network to estimate missing pixels. E4S (liu2023fine, ) redefines the face-swapping problem as a mask-exchanging problem for specific information. It utilizes a mask-guided injection module to perform face swapping in the latent space of StyleGAN. However, overall, the methods above have not thoroughly addressed the issue of artifacts in generated images under extreme occlusion conditions.
@@
6) Trade-offs between Identity-replacement and Attribute-retention. In addition to the occlusion issues that need further handling, researchers (gao2021information, ; shiohara2023blendface, ) discover that the balance between identity replacement and attribute preservation in generated images seems akin to a seesaw. Many methods (xu2022styleswap, ; kim2022smooth, ; liu2023high, ) explore the equilibrium between identity replacement and attribute retention. StyleSwap (xu2022styleswap, ) introduces a novel swapping guidance strategy, the ID reversal, to enhance the similarity of facial identity in the output. Shiohara et al(shiohara2023blendface, ) propose BlendFace, using an identity encoder that extracts identity features from the source image and uses it as identity distance loss, guiding the generator to produce facial exchange results.
@@
7) Model Light-weighting is also an important topic with profound implications for the widespread application of models. FastSwap (yoo2023fastswap, ) achieves this by innovating a decoder block called Triple Adaptive Normalization (TAN), effectively integrating identity information from the source image and pose information from the target image. XimSwap (ancilotto2023ximswap, ) modifies the design of convolutional blocks and the identity injection mechanism, successfully deploying on STM32H743.
$\bullet$ Diffusion-based. The latest studies (kim2025diffface, ; zhao2023diffswap, ; han2023generalist, ; liu2024towards, ) in this area produce promising generation results. DiffSwaps (zhao2023diffswap, ) redefines the face swapping problem as a conditional inpainting task. Liu et al(liu2024towards, ) introduce a multi-modal face generation framework and achieved this by introducing components such as balanced identity and expression encoders to the conditional diffusion model, striking a balance between identity replacement and attribute preservation during the generation process. As a novel facial generalist model, FaceX (han2023generalist, ) can achieve various facial tasks, including face swapping and editing. Leveraging the pre-trained StableDiffusion (blattmann2023stable, ) has significantly improved the quality and model training speed. Baliah et al(baliah2025realistic, ) achieves improved identity fidelity, pose consistency, and generalization through self-supervised inpainting, DDIM multi-step sampling, CLIP-based feature disentanglement, and mask shuffling.
$\bullet$ Alternative Techniques. Some methods stand independently from the above classifications that are discussed here collectively. Fast Face-swap (korshunova2017fast, ) views the identity swap task as a style transfer task, achieving its goals based on VGG-Net. However, this method has poor generalization. Some methods (cui2023face, ; cao2023transfs, ) apply the Transformer architecture to face swapping tasks. Leveraging a facial encoder based on the Swin Transformer (liu2021swin, ), TransFS (cao2023transfs, ) obtains rich facial features, enabling facial swapping in high-resolution images. CanonSwap (luo2025canonswap, ) introduces a motion–appearance disentangled video face swapping framework that enables accurate identity transfer while preserving fine-grained temporal dynamics.
@@
Table 3. Overview of face reenactment methods. Notations: ➊ Voxceleb, ➋ Self-build, ➌ Voxceleb2, ➍ TalkingHead-1KH, ➎ CelebV-HQ, ➏ VFHQ, ➐ RaFD, ➑ VGGFace, ➒ CelebV, ➓FFHQ. Expression (Exp).
@@
Kim et al(kim2018deep, ) TOG’18 Exp, Pose, Blink Using synthesized rendering images of a parameterized face model as input, creating lifelike video frames for the target actor.
Kim et al(kim2019neural, ) TOG’19 Lip,Exp, Pose Built on a recurrent generative adversarial network, it employs a hierarchical neural face renderer to synthesize realistic video frames.
HeadGAN (doukas2021headgan, ) ECCV’21 Exp, Pose Using 3DMM for facial modeling provides a 3D prior to the GAN, effectively guiding the generator to accurately recover pose and expression from the target frame.
Face2Faceρ (yang2022face2face, ) ECCV’22 Exp, Pose Decoupling the actor’s facial appearance and motion information with two separate encodings allows the network to learn facial app -earance and motion priors.
PECHead (gao2023high, ) CVPR’23 Lip, Exp, Pose ➌➍➎➏ A novel multi-scale feature alignment module for motion perception is proposed to minimize distortion during motion transmission.
Maskrenderer (behrouzi2025maskrenderer, ) PR’25 Lip, Exp, Pose The framework is robust to occlusion and large mismatches between Source and Driving facial structures.
@@
Zakharov et al(zakharov2019few, ) ICCV’19 Exp, Pose ➊➌ Proposed a meta-learning framework for adversarial generative models, reducing the required training data size.
FReeNet (zhang2020freenet, ) CVPR’20 Exp, Pose (moore2010multi, ) A new triple perceptual loss is proposed to richly reproduce facial details of the face.
DG (hsu2022dual, ) CVPR’22 Exp, Pose ➊➌➐ A proposed dual generator model network for large pose face reproduction.
Doukas et al(doukas2023free, ) TPAMI’23 Exp, Pose, Gaze (kellnhofer2019gaze360, )(zhang2015appearance, ) Eye gaze control in the generated video is implemented to further enhance visual realism.
MetaPortrait (zhang2023metaportrait, ) CVPR’23 Exp, Pose By establishing dense facial keypoint matching, accurate deformation field prediction is achieved, and the model training is expedited based on the meta-learning philosophy.
Yang et al(yang2023learning, ) AAAI’24 Exp, Pose ➊➌➍ The facial tri-plane is represented by canonical tri-plane, identity deformation, and motion components, achieving face reenactment without the need for 3D parameter model priors.
FSRT(rochow2024fsrt, ) CVPR’24 Exp, Pose The Transformer-based encoder-decoder effectively encodes attributes and improves action transmission quality.
DiffusionAct(bounareli2025diffusionact, ) FG’25 Exp, Pose The nethod allows one-shot, self, and cross-subject reenactment, without requiring subject-specific fine-tuning.
@@
HiDe-NeRF (li2023one, ) CVPR’23 Lip, Exp, Pose ➊➌➍ High-fidelity and free-viewing talking head synthesis using deformable neural radiation fields.
HyperReenact (bounareli2023hyperreenact, ) ICCV’23 Exp, Pose ➊➌ Exploiting the effectiveness of hypernetworks in real image inversion tasks and extending them to real image manipulation.
Stylemask (bounareli2023stylemask, ) FG’23 Exp, Pose This work optimizes a Mask Network and combines it with StyleGAN2’s style potential space S in order to achieve the separation of facial pose and expression of the target image from the identity features of the source image.
Bounareli et al(bounareli2024one, ) IJCV’24 Exp, Pose ➊➌ In GANs’ latent space, head pose and expression changes are decoupled, achieving near-real outputs through real image embedding.
Chang et al(chang2025enhancing, ) AAAI’25 Exp, Pose ➎ (zhang2021flow, ) Introduce a StyleGAN2-based end-to-end framework for high-fidelity one-shot video reenactment at 1024 resolution, with conditional disentanglement and feature-space refinement improving accuracy and fine-detail preservation
@@
ICface (tripathy2020icface, ) WACV’20 Exp, Pose The model is decoupled and driven by interpretable control signals that can be obtained from multiple sources such as external driving videos and manual controls.
@@
Oorloff et al(oorloff2023robust, ) ICCV’23 Lip, Exp, Pose Identity and attribute decomposition are realized in StyleGAN2’s latent space, and a cyclic manifold adjustment technique enhances facial reconstruction results.
@@
Xue et al. (xue2023high, ) TOMM’23 Exp, Pose ➊➌ High-fidelity facial generation is achieved by using information-rich Projected Normalized Coordinate Code (PNCC) and eye maps, replacing sparse facial landmark representations.
3.1.2. Face Reenactment